Casey and David leave school at the same time and drive in opposite directions. Casey travels at a speed that is 15 mph faster than David. After 3 hours, they are 261 miles apart. Find the rateof speed for both Casey and David
Answer by mananth(16949)

David x mph
Casey x+ 15 mph
..
They are moving away from each other

So add up their speed.

combined speed = x+ 1 x + 15
( 2 x + 15 )
Time = 3 hours
Distance = 261 miles
Distance = speed * time
( 2 x 15 )* 3 = 261
6 x + 45 = 261
6 x = 261 + -45
6 x = 216
/ 6
x= 36 mph David
Casey 36 + 15 = 51 mph